How to calculate degree of hardness of water with an example?

Grade:12

1 Answers

Vikas TU
4 years ago
Dear student 
degree of hardness=strength of hardness producing substance × multiplication factor 
so,multiplication factor=100/2 × chemical equivalent of hardness producing substance.
Example 
we need to calculate the hardness if 9.6 mg of Mgso4 present in 1 kg of water , 
so, here 
given , 1000g of water contains Mgso4 = 9.6 mg 
1000000g of water contain mgso4 = 9.6 * 1000 = 9600 mg = 9.6 g 
u need to remember – 100000g = one million parts of water 
1 mole of mgso4 = 1 mole of caco3 
120 g of mgso4 = 100g of caco3 
9.6 g of mgso4 = (100g * 9.6 g ) / 120 g of caco3 = 8g of caco3 
so, here degree of hardness = 8 ppm